OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x^2 - x - 30 A. x + 3 B. x - 5 C. x - 6 D. x + 10

OpenStudy (cwrw238):

let x^ 2 - x - 30 = ( x + a)(x + b) ab must be = -30 and a + b = -1 this should help

OpenStudy (anonymous):

so what letter is it @cwrw238

OpenStudy (anonymous):

C

OpenStudy (cwrw238):

dont just give the answer Dtan1293... plz the read open study guidelines

OpenStudy (anonymous):

oh my bad, won't happen again

OpenStudy (cwrw238):

5*-6 = -30 and -6 + 5 = -1 so we get (x + 5)( x - 6) C is correct